Figure 1: From Yvette Guilbert’s How to Sing a Song

Number and Description  PU AR TS
1 Pleasant surprise  8.1 8.1 7.2
2 On plane – growing anxiety*  3.8 6.8 4.5
3 Savoring Coca-Cola (“candid” shot)  4.9 4.6 4.7
4 Anticipation of refreshing drink  6.4 6.6 5.8
5 On plane – increasing anxiety*  3.0 4.6 3.7
6 On plane which is starting to fall*  1.3 8.6 8.6
[7] Grief – mother has just died of cancer  3.8 4.2 4.1
8 Strain – pulling at finger tips  6.6 6.7 6.6
9 Extreme pain  3.9 3.0 3.3
10 Anger at dog beating  1.3 8.8 8.9
11 Intense anger – during heated argument  2.2 7.1 7.7
12 Disdain  4.4 3.8 3.1
[13] Savoring drink (Coca-Cola)  5.9 5.4 4.8
14 Repose after anger  6.9 6.0 5.0
[15] Very pleasant surprise (told she is  8.8 7.8 7.1

going to get lead in play)

[16] Maternal love – newly adopted baby in arms 7.0 5.9 4.0
17 Pleasant surprise  5.7 4.7 4.5
18 laughter at joke (“candid” shot)  8.5 7.9 7.4
19 Very pleasant surprise  8.7 8.9 8.9
[20] Physical exhaustion  3.3 2.5 3.1
21 Mild anger – being kept waiting  3.7 3.8 3.4
22 Intense anger – argument with play director 2.2 5.8 7.1
24 Alerted – senses something wrong  4.3 4.6 4.9

with plane she is in*

26 Watching her baby  6.1 5.8 5.4
[28] Beginning of anxiety – see  3.5 6.1 6.8

smoking plane engine*

[29] Anger at seeing dog beaten  2.1 8.0 8.2
[30] Physical strain – pulling very hard  6.7 4.2 6.6

on seat of chair

[32] Unexpected meeting with old boy friend  7.4 6.8 5.9

of whom she is still fond

33 Mild physical discomfort  4.5 3.4 3.2
35 Response to amusing story  8.2 7.6 6.9
[36] Revulsion – seeing the face of a  2.9 3.0 5.1

woman covered with sores

[37] Extreme pain – lancing boil  2.2 2.2 6.4

38 Chiding boy friend  5.9 5.7 4.6
39 Photographer teasing  8.8 8.0 7.8
41 Revulsion – seeing a cat mangled 3.6 3.3 4.2

by a car

42 Looking up at lover  6.1 6.2 5.1
43 Mild anger  6.3 6.1 5.0
45 Beginning of anxiety*  4.1 4.2 3.4
46 Receiving posthumous war medal 4.9 4.8 4.2

awarded to husband

47 Strong anger – at dog beaten  2.6 6.3 7.2
48 Pleasant surprise  6.9 6.8 5.8
49 Physical exhaustion  4.0 3.0 2.2
50 Desire to adopt child  6.3 6.1 4.9
[51] Knowledge that plane she is on 1.1 8.6 8.9

will crash any moment*

52 Disgust  2.3 2.4 4.2
54 Pain – having boil lanced  3.8 2.9 3.6
55 Very pleasant surprise – told she 8.1 8.1 7.7

will be on magazine cover

[56] Light sleep on airplane*  4.1 1.3 1.0

These 48 pictures were rated by 96 students in the introductory course at Brown University. The pictures marked with an asterisk (*) were taken during an imagined sequence which started with the girl dozing in a plane (56), continued as she was alerted by a noise (45), showed her in a state of gradually building fear as she realized the danger of a crash (5,2,24,6,28), and ended as she screamed (51) just as the plane crashed.

The numbers in brackets are the faces used in the Abelson-Sermat experiment, and are numbered in (the same) order from 1 to 13.

PU: Pleasant/Unpleasant; AR: Attention/Rejection; TS: Tension/Sleep

Table 1: Proximities (Dissimilarities) for Faces

:

1  2  3  4  5  6  7  8  9  10 11  12  13
1 0.00 4.05 8.25 5.57 1.15 2.97 4.34 4.90 6.25 1.55 1.68 6.57 3.93
2 4.05 0.00 2.54 2.69 2.67 3.88 8.53 1.31 1.88 4.84 5.81 7.43 4.51
3 8.25 2.54 0.00 2.11 8.98 9.27 11.87 2.56 0.74 9.25 7.92 8.30 8.47
4 5.57 2.69 2.11 0.00 3.78 6.05 9.78 4.21 0.45 4.92 5.42 8.93 3.48
5 1.15 2.67 8.98 3.78 0.00 2.34 7.12 5.90 4.77 2.22 4.34 8.16 1.60
6 2.97 3.88 9.27 6.05 2.34 0.00 1.36 5.18 5.45 4.17 4.72 4.66 4.89
7 4.34 8.53 11.87 9.78 7.12 1.36 0.00 8.47 10.20 5.44 4.31 1.57 9.18
8 4.90 1.31 2.56 4.21 5.90 5.18 8.47 0.00 2.63 5.45 3.79 6.49 6.05
9 6.25 1.88 0.74 0.45 4.77 5.45 10.20 2.63 0.00 7.10 6.58 9.77 6.55
10 1.55 4.84 9.25 4.92 2.22 4.17 5.44 5.45 7.10 0.00 1.98 4.93 4.12
11 1.68 5.81 7.92 5.42 4.34 4.72 4.31 3.79 6.58 1.98 0.00 4.83 3.51
12 6.57 7.43 8.30 8.93 8.16 4.66 1.57 6.49 9.77 4.93 4.83 0.00 12.65
13 3.93 4.51 8.47 3.48 1.60 4.89 9.18 6.05 6.55 4.12 3.51 12.65 0.00

Original data from Abelson & Sermat, JEP, 1962, 546–554 — based on (transformed) averages over 30 subjects (originally) judging all pairs on a 9-point scale as to a difference in emotional expression or content.

Legend: 1: [7] Grief at death of mother 2: [13] Savoring a coke 3: [15] Very pleasant surprise 4: [16] Maternal love — baby in arms 5: [20] Physical exhaustion 6: [23] Something wrong with plane 7: [29] Anger seeing dog beaten 8: [30] Pulling hard on seat of chair 9: [32] Unexpectedly meets old boyfriend 10: [36] Revulsion 11: [37] Extreme pain 12: [51] Knows plane will crash 13: [56] Light sleep

Figure 2: City-Block Scaling

The coordinates for the embedded properties are indicated by x’s: (0,0) and (.69,.31) for PU; (-.10,1.06) for AR; (-.30,1.29) for TS.

Squared correlations (variances-accounted-for) as lines embedded in the City-Block configuration: PU (90.0%); AR (42.2%); TS (90.0%).

Axis one correlation to PU (.93); AR (-.11); TS (-.40); axis two correlation to PU (.23); AR (.64); TS (.84)

1: [7] Grief at death of mother; 2: [13] Savoring a coke; 3: [15] Very pleasant surprise; 4: [16] Maternal love — baby in arms; 5: [20] Physical exhaustion; 6: [23] Something wrong with plane; 7: [29] Anger seeing dog beaten; 8: [30] Pulling hard on seat of chair; 9: [32] Unexpectedly meets old boyfriend; 10: [36] Revulsion; 11: [37] Extreme pain; 12: [51] Knows plane will crash; 13: [56] Light sleep

Figure 3: Euclidean Scaling

The coordinates for the embedded properties are indicated by x’s: (0,0) and (.65,-.28) for PU; (-.06,-.89) for AR; (-.27,-.80) for TS.

Squared correlations (variances-accounted-for) as lines embedded in the City-Block configuration: PU (94.1%); AR (75.7%); TS (90.0%).

Axis one correlation to PU (-.93); AR (.09); TS (.42); axis two correlation to PU (-.27); AR (-.86); TS (-.85)

1: [7] Grief at death of mother; 2: [13] Savoring a coke; 3: [15] Very pleasant surprise; 4: [16] Maternal love — baby in arms; 5: [20] Physical exhaustion; 6: [23] Something wrong with plane; 7: [29] Anger seeing dog beaten; 8: [30] Pulling hard on seat of chair; 9: [32] Unexpectedly meets old boyfriend; 10: [36] Revulsion; 11: [37] Extreme pain; 12: [51] Knows plane will crash; 13: [56] Light sleep

Figure 4: Comparison of City-Block and Euclidean Scaling

City-Block coordinates are ‘x’; Euclidean coordinates are ‘.’; transformed Euclidean coordinates to the City-Block target are ‘o’.

1: [7] Grief at death of mother; 2: [13] Savoring a coke; 3: [15] Very pleasant surprise; 4: [16] Maternal love — baby in arms; 5: [20] Physical exhaustion; 6: [23] Something wrong with plane; 7: [29] Anger seeing dog beaten; 8: [30] Pulling hard on seat of chair; 9: [32] Unexpectedly meets old boyfriend; 10: [36] Revulsion; 11: [37] Extreme pain; 12: [51] Knows plane will crash; 13: [56] Light sleep

Partition Hierarchy  Level Formed

{9,4,3,2,8,13,5,11,10,1,6,12,7}  7.42
{9,4,3,2,8,13,5,11,10,1,6},{12,7}  5.84
{9,4,3,2,8},{13,5,11,10,1,6},{12,7}  3.82
{9,4,3,2,8},{13,5,11,10,1},{6},{12,7}  3.21
{9,4,3,2,8},{13,5},{11,10,1},{6},{12,7}  2.75
{9,4,3},{2,8},{13,5},{11,10,1},{6},{12,7}  1.83
{9,4,3},{2,8},{13,5},{11},{10,1},{6},{12,7}  1.60
{9,4,3},{2,8},{13},{5},{11},{10,1},{6},{12,7}  1.57
{9,4,3},{2,8},{13},{5},{11},{10,1},{6},{12},{7}  1.55
{9,4,3},{2,8},{13},{5},{11},{10},{1},{6},{12},{7}  1.43
{9,4},{3},{2,8},{13},{5},{11},{10},{1},{6},{12},{7}  1.31
{9,4},{3},{2},{8},{13},{5},{11},{10},{1},{6},{12},{7}  .45
{9},{4},{3},{2},{8},{13},{5},{11},{10},{1},{6},{12},{7}  —

Legend: 1: [7] Grief at death of mother 2: [13] Savoring a coke 3: [15] Very pleasant surprise 4: [16] Maternal love — baby in arms 5: [20] Physical exhaustion 6: [23] Something wrong with plane 7: [29] Anger seeing dog beaten 8: [30] Pulling hard on seat of chair 9: [32] Unexpectedly meets old boyfriend 10: [36] Revulsion 11: [37] Extreme pain 12: [51] Knows plane will crash 13: [56] Light sleep Table 2: Ultrametric fitted values (least-squares)[vaf of 54.6%]

9  4  3  2  8  13  5  11 10  1  6  12  7
9 .00 .45 1.43 2.75 2.75 5.84 5.84 5.84 5.84 5.84 5.84 7.42 7.42
4 .45 .00 1.43 2.75 2.75 5.84 5.84 5.84 5.84 5.84 5.84 7.42 7.42
3 1.43 1.43 .00 2.75 2.75 5.84 5.84 5.84 5.84 5.84 5.84 7.42 7.42
2 2.75 2.75 2.75 .00 1.31 5.84 5.84 5.84 5.84 5.84 5.84 7.42 7.42
8 2.75 2.75 2.75 1.31 .00 5.84 5.84 5.84 5.84 5.84 5.84 7.42 7.42
13 5.84 5.84 5.84 5.84 5.84 .00 1.60 3.21 3.21 3.21 3.82 7.42 7.42
5 5.84 5.84 5.84 5.84 5.84 1.60 .00 3.21 3.21 3.21 3.82 7.42 7.42
11 5.84 5.84 5.84 5.84 5.84 3.21 3.21 .00 1.83 1.83 3.82 7.42 7.42
10 5.84 5.84 5.84 5.84 5.84 3.21 3.21 1.83 .00 1.55 3.82 7.42 7.42
1 5.84 5.84 5.84 5.84 5.84 3.21 3.21 1.83 1.55 .00 3.82 7.42 7.42
6 5.84 5.84 5.84 5.84 5.84 3.82 3.82 3.82 3.82 3.82 .00 7.42 7.42
12 7.42 7.42 7.42 7.42 7.42 7.42 7.42 7.42 7.42 7.42 7.42 .00 1.57
7 7.42 7.42 7.42 7.42 7.42 7.42 7.42 7.42 7.42 7.42 7.42 1.57 .00

Note: for any three objects (say, 1, 2, and 3), the largest two proximities between the pairs are equal (p_{12} = 5.84; p_{13} = 5.84; p_{23} = 2.75)

Legend: 1: [7] Grief at death of mother 2: [13] Savoring a coke 3: [15] Very pleasant surprise 4: [16] Maternal love — baby in arms 5: [20] Physical exhaustion 6: [23] Something wrong with plane 7: [29] Anger seeing dog beaten 8: [30] Pulling hard on seat of chair 9: [32] Unexpectedly meets old boyfriend 10: [36] Revulsion 11: [37] Extreme pain 12: [51] Knows plane will crash 13: [56] Light sleep

Figure 5: A dendrogram representation for the ultrametric

1: [7] Grief at death of mother; 2: [13] Savoring a coke; 3: [15] Very pleasant surprise; 4: [16] Maternal love — baby in arms; 5: [20] Physical exhaustion; 6: [23] Something wrong with plane; 7: [29] Anger seeing dog beaten; 8: [30] Pulling hard on seat of chair; 9: [32] Unexpectedly meets old boyfriend; 10: [36] Revulsion; 11: [37] Extreme pain; 12: [51] Knows plane will crash; 13: [56] Light sleep

Table 3: Additive-tree fitted values (least-squares)[vaf of 82.3%] :

1  2  3  4  5  6  7  8  9  10 11  12  13
1 .00 4.06 7.48 5.22 3.03 2.47 5.34 4.60 5.85 1.55 1.61 5.53 4.47
2 4.06 .00 3.67 1.41 3.41 4.93 7.81 1.09 2.04 4.49 4.38 8.00 4.86
3 7.48 3.67 .00 2.09 6.83 8.34 11.22 4.50 .74 7.91 7.79 11.41 8.28
4 5.22 1.41 2.09 .00 4.57 6.08 8.96 2.24 .47 5.65 5.53 9.15 6.02
5 3.03 3.41 6.83 4.57 .00 3.89 6.77 3.95 5.20 3.46 3.35 6.96 1.60
6 2.47 4.93 8.34 6.08 3.89 .00 2.91 5.46 6.72 2.90 2.78 3.11 5.34
7 5.34 7.81 11.22 8.96 6.77 2.91 .00 8.34 9.59 5.77 5.66 1.57 8.22
8 4.60 1.09 4.50 2.24 3.95 5.46 8.34 .00 2.88 5.03 4.92 8.53 5.40
9 5.85 2.04 .74 .47 5.20 6.72 9.59 2.88 .00 6.28 6.17 9.79 6.65
10 1.55 4.49 7.91 5.65 3.46 2.90 5.77 5.03 6.28 .00 2.05 5.97 4.91
11 1.61 4.38 7.79 5.53 3.35 2.78 5.66 4.92 6.17 2.05 .00 5.85 4.79
12 5.53 8.00 11.41 9.15 6.96 3.11 1.57 8.53 9.79 5.97 5.85 .00 8.41
13 4.47 4.86 8.28 6.02 1.60 5.34 8.22 5.40 6.65 4.91 4.79 8.41 0.00

Note: for any four objects (say, 1, 2, 3, and 4), the largest two particular sums are equal (p_{12} + p_{34} = 4.06 + 2.09 = 6.15; p_{13} + p_{24} = 7.48 + 1.41 = 8.89; p_{14} + p_{23} = 5.22 + 3.67 = 8.89)

Legend: 1: [7] Grief at death of mother 2: [13] Savoring a coke 3: [15] Very pleasant surprise 4: [16] Maternal love — baby in arms 5: [20] Physical exhaustion 6: [23] Something wrong with plane 7: [29] Anger seeing dog beaten 8: [30] Pulling hard on seat of chair 9: [32] Unexpectedly meets old boyfriend 10: [36] Revulsion 11: [37] Extreme pain 12: [51] Knows plane will crash 13: [56] Light sleep Figure 6: An additive tree representation – all horizontal distances are considered zero and are used for display purposes only

Legend:

1: [7] Grief at death of mother; 2: [13] Savoring a coke; 3: [15] Very pleasant surprise; 4: [16] Maternal love — baby in arms; 5: [20] Physical exhaustion; 6: [23] Something wrong with plane; 7: [29] Anger seeing dog beaten; 8: [30] Pulling hard on seat of chair; 9: [32] Unexpectedly meets old boyfriend; 10: [36] Revulsion; 11: [37] Extreme pain; 12: [51] Knows plane will crash; 13: [56] Light sleep

Figure 7: An “angular” representation for the additive tree

Legend:

1: [7] Grief at death of mother; 2: [13] Savoring a coke; 3: [15] Very pleasant surprise; 4: [16] Maternal love — baby in arms; 5: [20] Physical exhaustion; 6: [23] Something wrong with plane; 7: [29] Anger seeing dog beaten; 8: [30] Pulling hard on seat of chair; 9: [32] Unexpectedly meets old boyfriend; 10: [36] Revulsion; 11: [37] Extreme pain; 12: [51] Knows plane will crash; 13: [56] Light sleep

Figure 8: A “radial” representation for the additive tree

Legend:

1: [7] Grief at death of mother; 2: [13] Savoring a coke; 3: [15] Very pleasant surprise; 4: [16] Maternal love — baby in arms; 5: [20] Physical exhaustion; 6: [23] Something wrong with plane; 7: [29] Anger seeing dog beaten; 8: [30] Pulling hard on seat of chair; 9: [32] Unexpectedly meets old boyfriend; 10: [36] Revulsion; 11: [37] Extreme pain; 12: [51] Knows plane will crash; 13: [56] Light sleep

Figure 9: The linear unidimensional scaling of the 13 faces

1: [7] Grief at death of mother; 2: [13] Savoring a coke; 3: [15] Very pleasant surprise; 4: [16] Maternal love — baby in arms; 5: [20] Physical exhaustion; 6: [23] Something wrong with plane; 7: [29] Anger seeing dog beaten; 8: [30] Pulling hard on seat of chair; 9: [32] Unexpectedly meets old boyfriend; 10: [36] Revulsion; 11: [37] Extreme pain; 12: [51] Knows plane will crash; 13: [56] Light sleep

Figure 10: The circular unidimensional scaling of the 13 faces

1: [7] Grief at death of mother; 2: [13] Savoring a coke; 3: [15] Very pleasant surprise; 4: [16] Maternal love — baby in arms; 5: [20] Physical exhaustion; 6: [23] Something wrong with plane; 7: [29] Anger seeing dog beaten; 8: [30] Pulling hard on seat of chair; 9: [32] Unexpectedly meets old boyfriend; 10: [36] Revulsion; 11: [37] Extreme pain; 12: [51] Knows plane will crash; 13: [56] Light sleep Table 4: Anti-Robinson fitted values (least-squares)[vaf of 91.1%]

12  7  6  11 10  1  5  13  2  8  4  9  3
12 .00 1.57 4.66 4.83 4.93 6.57 8.16 8.86 8.86 8.86 9.36 9.99 10.09
7 1.57 .00 1.36 4.31 4.89 4.89 7.12 8.73 8.73 8.73 9.36 9.99 10.09
6 4.66 1.36 .00 3.71 3.71 3.71 3.71 4.80 4.80 5.18 6.05 6.38 9.27
11 4.83 4.31 3.71 .00 1.83 1.83 3.71 3.85 4.80 4.80 5.42 6.38 8.60
10 4.93 4.89 3.71 1.83 .00 1.55 2.22 3.85 4.80 4.80 5.25 6.38 8.60
1 6.57 4.89 3.71 1.83 1.55 .00 1.15 3.85 4.05 4.80 5.25 6.25 8.60
5 8.16 7.12 3.71 3.71 2.22 1.15 .00 1.60 3.59 4.80 4.80 5.66 8.60
13 8.86 8.73 4.80 3.85 3.85 3.85 1.60 .00 3.59 4.77 4.77 5.66 8.47
2 8.86 8.73 4.80 4.80 4.80 4.05 3.59 3.59 .00 1.31 2.75 2.75 2.75
8 8.86 8.73 5.18 4.80 4.80 4.80 4.80 4.77 1.31 .00 2.75 2.75 2.75
4 9.36 9.36 6.05 5.42 5.25 5.25 4.80 4.77 2.75 2.75 .00 .45 2.11
9 9.99 9.99 6.38 6.38 6.38 6.25 5.66 5.66 2.75 2.75 .45 .00 .74
3 10.09 10.09 9.27 8.60 8.60 8.60 8.60 8.47 2.75 2.75 2.11 .74 .00

Legend: 1: [7] Grief at death of mother 2: [13] Savoring a coke 3: [15] Very pleasant surprise 4: [16] Maternal love — baby in arms 5: [20] Physical exhaustion 6: [23] Something wrong with plane 7: [29] Anger seeing dog beaten 8: [30] Pulling hard on seat of chair 9: [32] Unexpectedly meets old boyfriend 10: [36] Revulsion 11: [37] Extreme pain 12: [51] Knows plane will crash 13: [56] Light sleep